# Tail Distribution of Regret in Optimistic Reinforcement Learning

### v1

@@ Abstract
We derive instance-dependent tail bounds for the regret of optimism-based reinforcement learning in finite-horizon tabular Markov decision processes with unknown transition dynamics. Focusing on a UCBVI-type algorithm, we characterize the tail distribution of the cumulative regret $R_{K}$ over $K$ episodes, rather than only its expectation or a single high-probability quantile. We analyze two natural exploration-bonus schedules: (i) a $K$-dependent scheme that explicitly incorporates the total number of episodes $K$, and (ii) a $K$-independent scheme that depends only on the current episode index. For both settings, we obtain an upper bound on $\Pr(R_{K}\geq x)$ that exhibits a distinctive two-regime structure: a sub-Gaussian tail starting from an instance-dependent scale $m_{K}$ up to a transition threshold, followed by a sub-Weibull tail beyond that point. We further derive corresponding instance-dependent bounds on the expected regret $\mathbb{E}[R_{K}]$. The proposed algorithm depends on a tuning parameter $\alpha$, which balances the expected regret and the range over which the regret exhibits a sub-Gaussian tail. To the best of our knowledge, our results provide one of the first comprehensive tail-regret guarantees for a standard optimistic algorithm in episodic reinforcement learning.
@@ 1  Introduction
We consider online Reinforcement Learning (RL) in finite-horizon Markov decision processes (MDPs) with unknown transition dynamics. In this setting, a learning agent interacts with an episodic environment for $K$ episodes of length $H$, and aims to minimize its cumulative regret with respect to an optimal policy. A long line of work has established sharp worst-case regret guarantees for model-based and model-free algorithms in tabular MDPs; see, for instance, the near-minimax optimality results for UCRL2 under the average-reward criterion (jaksch2010near) and for UCBVI and its variants in the episodic setting (azar2017minimax; zanette2019tighter). These results show that, up to logarithmic factors, the optimal worst-case regret scales as $\widetilde{\mathcal{O}}(\sqrt{HSAK})$, where $S$ and $A$ are the sizes of the state and action spaces.
More recently, there has been substantial progress on instance-dependent analysis in RL. Inspired by the classical gap-dependent theory of stochastic bandits, several works have derived regret bounds that depend on the suboptimality gaps of state–action pairs rather than only on $(S,A,H,K)$. For tabular MDPs, this includes non-asymptotic gap-dependent bounds for model-based algorithms (simchowitz2019non), fine-grained gap-dependent bounds via adaptive multi-step bootstrapping (xu2021fine), and gap-dependent guarantees for model-free $Q$-learning and its refinements (yang2020q; velegkas2022reinforcement; zheng2025gap). These results provide a more refined characterization of the regret in MDPs by incorporating an instance-dependent $Q^{*}$-gap, which measures the separation between optimal and suboptimal actions at each state.
@@
for a fixed confidence level $\delta\in(0,1)$, typically by combining a union bound with martingale concentration inequalities. While such high-probability guarantees are useful, they only describe the regret distribution at a single point and require the user to choose $\delta$ in advance. By contrast, in many safety-critical applications, such as healthcare decision making, robotic control in human environments, and financial risk management inamdar2024comprehensive, one would like a more detailed characterization of the entire tail distribution of regret, that is, explicit bounds on $\mathbb{P}(R_{K}\geq x)$ that hold for all thresholds $x\geq\mathbb{E}[R_{K}]$.
In this paper, we take a step towards closing this gap by studying the concentration of regret for a UCBVI-type algorithm in finite-horizon MDPs. We show that the tail of the cumulative regret, $P(R_{K}\geq x)$, admits an explicit, non-asymptotic sub-Gaussian or sub-Weibull bound, depending on the value of $x$ and the instance-dependent $Q^{\star}$-gap parameter.
@@
  • For the $K$-dependent scheme, we show that the regret exhibits a sub-Gaussian tail on an intermediate regime of $x$ and a sub-Weibull tail beyond $x>\tilde{\mathcal{O}}\!\big(H^{3/2}K^{(1+\alpha)/2}\big)$, where $\alpha\in[0,1]$ is a tuning parameter.
  • For the $K$-independent scheme, we obtain a similar two-regime behavior, where the transition to sub-Weibull happens at $\tilde{\mathcal{O}}\!\big(H^{3/(2-\alpha)}K^{1/(2-\alpha)}\big)$.
  • We further show that the expected regret is $\mathbb{E}[R_{K}]\;\leq\;\tilde{\mathcal{O}}\!\Big(H^{4}SA(S+K^{\alpha})/\texttt{gap}^{*}\Big)$, where $\texttt{gap}^{*}>0$ is the global $Q^{\star}$-gap. The tuning parameter $\alpha$ balances the expected regret and the range over which the regret exhibits a sub-Gaussian tail.
@@ 2  Related Work
The modern theory of online RL in finite MDPs starts from worst-case regret guarantees under the optimism in the face of uncertainty principle. In the average-reward setting, the UCRL2 algorithm of jaksch2010near achieves near-minimax regret of order $\widetilde{\mathcal{O}}\!\big(DS\sqrt{AK}\big)$, where $D$ is the diameter of the MDP, and lower bound $\widetilde{\mathcal{O}}\!\big(\sqrt{DSAK}\big)$ show that the dependence on $(D,S,A,K)$ is almost optimal. Subsequent refinements sharpened the constants and logarithmic factors, for instance by tightening the confidence sets around the transition kernel via empirical Bernstein-type bonuses (bourel2020tightening; fruit2020improved).
In the episodic finite-horizon setting considered in this paper, the optimal worst-case scaling was established by the UCBVI algorithm of azar2017minimax, which attains regret $\widetilde{\mathcal{O}}\!\big(H\sqrt{SAK}+H\sqrt{KH}\big)$ and derive information-theoretic lower bounds of $\widetilde{\mathcal{O}}\!\big(H\sqrt{SAK}\big)$. A number of works have since refined or extended this result: efroni2019tight proposed the EULER algorithm, which yields worst-case bounds with improved dependence on the horizon; see also the recent minimax analyses and extensions to factored MDPs, e.g., tian2020towards. These results, however, are primarily worst-case and focus on bounding either the expectation or a single high-probability quantile of the regret, without characterizing the entire tail behavior.
@@
Motivated by the classical gap-dependent theory for stochastic bandits (see, e.g., the survey of bubeck2012regret), there has been a surge of interest in instance-dependent regret bounds for tabular MDPs. simchowitz2019non derived the first non-asymptotic gap-dependent regret bounds for model-based tabular RL, showing that the regret can scale polylogarithmically in $K$ (up to polynomial factors in $(S,A,H)$ and $1/\texttt{gap}^{\star}$). zanette2019tighter introduces the environmental norm, a variance-like quantity that controls the difficulty of exploration, and designed an optimistic algorithm whose regret depends on this norm rather than only on worst-case parameters. More recently, dann2021beyond developed improved instance-dependent bounds that replace the global value-function gaps with more local notions tailored to the subset of states visited by optimal policies, together with new information-theoretic lower bounds showing the necessity of these refined gap notions. On the model-free side, yang2020q showed that optimistic $Q$-learning enjoys logarithmic regret in episodic tabular MDPs under a global suboptimality-gap on $Q^{\star}$, thus establishing that simple temporal-difference methods can match the best-possible gap-dependent rates of model-based algorithms. xu2021fine proposed an adaptive multi-step bootstrapping scheme that yields tighter instance-dependent bounds.
Under the linear MDP assumption with function approximation, he2021logarithmic establish instance-dependent logarithmic regret for LSVI-UCB and UCRL-VTR in linear and linear-mixture MDPs. velegkas2022reinforcement further refine this picture by designing algorithms that incur only a bounded number of policy switches. Most recently, zhao2025logarithmic extend these guarantees to adversarial environments.
@@
A recent line of work, starting from fan2024fragility, shows that bandit algorithms that are information-theoretically optimized for Lai–Robbins optimal expected regret can exhibit extremely heavy-tailed regret distributions: the regret tail behaves like a truncated Cauchy, and higher moments grow polynomially in the horizon. Building on this observation, simchi2022simple; simchi2023stochastic design policies that explicitly trade off optimality and tail risk, achieving worst-case optimal (or near-optimal) expected regret while simultaneously enforcing exponentially light tails for large-regret events and characterizing sharp Pareto frontiers between expected regret and tail probabilities.
In contrast, RL work that explicitly targets tail or risk-sensitive criteria usually changes the performance objective rather than analyzing the tail of classical regret. For example, there is a growing literature on risk-sensitive or distributional RL that optimizes mean-variance, CVaR, or more general optimized certainty equivalent (OCE) criteria in MDPs; see, e.g., tamar2015optimizing; chow2015risk; xu2023regret and references therein. These works derive sample-complexity or regret guarantees for the risk-sensitive value of a policy, not for the tail of the standard regret with respect to the risk-neutral optimal policy.
@@ 3  Problem Formulation
$\texttt{gap}^{*}\;\coloneqq\;\min_{\begin{subarray}{c}(s,h)\in\mathcal{S}\times[H]\\ a\notin\mathcal{A}_{h}^{*}(s)\end{subarray}}\Big(Q_{h}^{*}(s,\pi_{h}^{*}(s))-Q_{h}^{*}(s,a)\Big),$ (1)
where $\pi^{*}\in\mathbf{\Pi}^{*}$ is any optimal Markov policy.
@@ 4  Main Results
We analyze a UCBVI-type algorithm that uses the following bonus function:
@@
Here $k\in\{0,\dots,K-1\}$ denotes the episode index, $n\geq 1$ denotes the number of visits to the corresponding state–action pair, KD refers to the bonus used in the $K$-dependent algorithm, and KI refers to the bonus used in the $K$-independent algorithm.
The term $K^{\alpha}$ in the definition of $b_{h}^{k}(\cdot)$ plays the role of a global exploration budget that directly controls the aggressiveness of the bonuses. In the original minimax UCBVI analysis of azar2017minimax, as well as in the subsequent instance-dependent analysis of simchowitz2019non, the exploration terms scale only logarithmically with the time horizon, roughly as $\sqrt{{\log(KSAH)}/{n}}$, which is tailored to control of the expected regret. In contrast, our bonuses scale as $\sqrt{(S\ln 2+\mu K^{\alpha})/n}$ (or $\sqrt{(S\ln 2+\mu(k+1)^{\alpha})/n}$ in the $K$-independent case), so that the uncertainty radius around the empirical model grows polynomially in $K$. This more conservative choice incentivizes sustained exploration even in later episodes: when the algorithm has visited a state–action pair only moderately many times, the bonus remains substantial and pushes the optimistic value estimates upward. From a tail perspective, this is precisely what allows us to derive nontrivial bounds on $\mathbb{P}(R_{K}\geq x)$ for large thresholds $x$, at the price of a controlled polynomial dependence on $K$ inside the bonus which will affect the expected regret.
Algorithm 1 UCBVI (Episodic Value Iteration with Stagewise UCB Bonus)
1:Input: $H,\mathcal{S},\mathcal{A}$, rewards $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$, initial $s_{0}$, bonus schedules $\{b^{k}_{h}(\cdot)\}$.
2:Init: For all $(s,a,h)$, $n_{h}^{0}(s,a)=0$ and $\hat{P}_{h}^{0}(\cdot\mid s,a)=\tfrac{1}{S}$.
@@
$a_{h}^{k}=\pi_{h}^{k}(s_{h}^{k})\quad\text{and}\quad s_{h+1}^{k}\sim P_{h}^{*}(\cdot\mid s_{h}^{k},a_{h}^{k}),\qquad\forall h\in[H]$
6:  Update: For all $(s,a,h)$,
@@
Fix $\alpha\in[0,1]$. For all $\gamma\in[0,1]$, the tail of the cumulative regret $R_{K}$ satisfies
$\mathbb{P}(R_{K}\geq x)\ \leq\ \exp\!\left(-\frac{\big(x-H\lceil K^{\gamma}\rceil-\,m_{K}\big)_{+}^{\,2}}{2H^{3}K}\right)\;+\;\delta_{K},$
@@
$m_{K}\;\coloneqq\frac{4H^{2}(H-1)(2H-1)SA({S\ln 2+2\mu K^{\alpha}})}{3\texttt{gap}^{*}},$ (2)
@@
$\delta_{K}=SAH\begin{cases}\displaystyle K\exp\left(-2\mu K^{\alpha}\right),&\text{{KD}},\\[8.61108pt] \displaystyle K\exp\left(-2\mu K^{\gamma\alpha}\right),&\text{{KI}}.\end{cases}$ (3)
Theorem 1 shows that, under a global $Q^{\star}$-gap, as defined in (1), the tail of the cumulative regret $R_{K}$ admits a decomposition into a sub-Gaussian component and a residual term $\delta_{K}$. The quantity $m_{K}$ is the baseline instance-dependent level of regret around which the distribution concentrates: it aggregates the bonuses incurred before all suboptimal actions have been sufficiently explored and eliminated, and it is of order
$m_{K}\;\approx\;\mathcal{O}\!\big(H^{4}SA(S+K^{\alpha})/\texttt{gap}^{*}\big)$
up to logarithmic factors. For thresholds $x$ that are moderately larger than $m_{K}$, the bound
$\mathbb{P}(R_{K}\geq x)\;\lesssim\;\exp\!\left(-\frac{(x-H\lceil K^{\gamma}\rceil-m_{K})_{+}^{\,2}}{2H^{3}K}\right)\;+\;\delta_{K}$
is of sub-Gaussian form with variance proxy $H^{3}K$, reflecting the martingale structure of the regret fluctuations after the main exploration phase.
Theorem 2.
@@
$\mathbb{E}[R_{K}]\ \leq\ \,m_{K}\;+H\lceil K^{\gamma}\rceil+\;2H^{2}\,(K-\lceil K^{\gamma}\rceil)\,\delta_{K},$
where $m_{K}$ is given in (2) and $\delta_{K}$ in (3).
Theorem 2 obtain an instance-dependent upper bound on the expected regret. The first two terms, $m_{K}+H\lceil K^{\gamma}\rceil$, capture the contribution of the baseline instance-dependent level of regret and the initial “burn-in” phase. The final term, proportional to $(K-\lceil K^{\gamma}\rceil)\delta_{K}$, accounts for the rare events in which the empirical model falls outside the “good” set at some episodes. In such events, the estimated transition kernel $\hat{\bm{P}}^{k}$ deviates too far from the true kernel $\bm{P}^{*}$ for some $k\in[K]$, causing the optimism condition to fail. Thanks to the exponential decay of $\delta_{K}$ in $K^{\alpha}$ (or $K^{\gamma\alpha}$), this residual contribution is negligible: for any fixed $\alpha>0$ and suitable choice of the constant $\mu>0$, the product $K\delta_{K}$ vanishes.
The tuning parameter $\alpha\in[0,1]$ controls the tradeoff between the expected regret and the decay rate of the residual term $\delta_{K}$. Larger values of $\alpha$ make the bonuses more aggressive, which increases the upper bound on the $\mathbb{E}[R_{K}]$ but also causes $\delta_{K}$ to decay faster with $K$ (either as $\exp(-2\mu K^{\alpha})$ or as $\exp(-2\mu K^{\gamma\alpha})$ depending on the bonus schedule). In this sense, $\alpha$ interpolates between prioritizing a smaller expected regret ($\alpha$ small) and enforcing lighter tails for very large deviations ($\alpha$ closer to one). By contrast, the parameter $\gamma\in[0,1]$ is not a tuning knob of the algorithm: it appears only in the analysis and controls the length of the initial burn-in phase $K^{\gamma}$ during which we make no attempt to exploit tail concentration. In Corollary 1, we choose $\gamma$ as a function of $x$ to simplify the resulting tail bound into a clean two-regime description.
@@
$\mathbb{P}(R_{K}\geq x)\leq\begin{cases}\exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{2}}{H^{3}K}\right)\right)&\text{if }\mathcal{O}(H^{4}SA(S+K^{\alpha})/\texttt{gap}^{*})\leq x\leq\tilde{\mathcal{O}}(H^{3/2}K^{(1+\alpha)/2})\\ \exp\left(-\tilde{\mathcal{O}}\left(K^{\alpha}\right)\right)&\text{if }x\geq\tilde{\mathcal{O}}(H^{3/2}K^{(1+\alpha)/2})\end{cases}$
@@
$\mathbb{P}(R_{K}\geq x)\leq\begin{cases}\exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{2}}{H^{3}K}\right)\right)&\text{if }\mathcal{O}(H^{4}SA(S+K^{\alpha})/\texttt{gap}^{*})\leq x\leq\tilde{\mathcal{O}}(H^{3/(2-\alpha)}K^{1/(2-\alpha)})\\ \exp\left(-\tilde{\mathcal{O}}\left(x^{\alpha}\right)\right)&\text{if }x\geq\tilde{\mathcal{O}}(H^{3/(2-\alpha)}K^{1/(2-\alpha)})\end{cases}$
Furthermore, $\mathbbm{E}[R_{K}]\leq\mathcal{O}(H^{4}SA(S+K^{\alpha})/\texttt{gap}^{*})$.

### v2

@@ Abstract
We derive instance-dependent tail bounds for the regret of optimism-based reinforcement learning in finite-horizon tabular Markov decision processes with unknown transition dynamics. Focusing on a UCBVI-type algorithm, we characterize the tail distribution of the cumulative regret $R_{K}$ over $K$ episodes, rather than only its expectation or a single high-probability quantile. We analyze two natural exploration-bonus schedules: (i) a $K$-dependent scheme that explicitly incorporates the total number of episodes $K$, and (ii) a $K$-independent scheme that depends only on the current episode index. For both settings, we obtain an upper bound on $\mathbb{P}(R_{K}\geq x)$ that exhibits a distinctive two-regime structure: a sub-Gaussian tail starting from an instance-dependent scale $m_{K}^{\mathrm{mb}}$ up to a transition threshold, followed by a sub-Weibull tail beyond that point. We further derive corresponding instance-dependent bounds on the expected regret $\mathbb{E}[R_{K}]$. The proposed algorithm depends on a tuning parameter $\alpha$, which balances the expected regret and the range over which the regret exhibits a sub-Gaussian tail. To the best of our knowledge, our results provide one of the first comprehensive tail-regret guarantees for a standard optimistic algorithm in episodic reinforcement learning.
@@ 1 Introduction
We consider online Reinforcement Learning (RL) in finite-horizon Markov decision processes (MDPs) with unknown transition dynamics. In this setting, a learning agent interacts with an episodic environment for $K$ episodes of length $H$, and aims to minimize its cumulative regret with respect to an optimal policy. A long line of work has established sharp worst-case regret guarantees for model-based and model-free algorithms in tabular MDPs; see, for instance, the near-minimax optimality results for UCRL2 under the average-reward criterion (Jaksch et al., 2010) and for UCBVI and its variants in the episodic setting (Azar et al., 2017; Zanette and Brunskill, 2019). These results show that, up to logarithmic factors, the optimal worst-case regret scales as $\widetilde{\mathcal{O}}(\sqrt{HSAK})$, where $S$ and $A$ are the sizes of the state and action spaces.
More recently, there has been substantial progress on instance-dependent analysis in RL. Inspired by the classical gap-dependent theory of stochastic bandits, several works have derived regret bounds that depend on the suboptimality gaps of state–action pairs rather than only on $(S,A,H,K)$. For tabular MDPs, this includes non-asymptotic gap-dependent bounds for model-based algorithms (Simchowitz and Jamieson, 2019), fine-grained gap-dependent bounds via adaptive multi-step bootstrapping (Xu et al., 2021), and gap-dependent guarantees for model-free $Q$-learning and its refinements (Yang et al., 2021; Velegkas et al., 2022; Zheng et al., 2024). These results provide a more refined characterization of the regret in MDPs by incorporating an instance-dependent $Q^{*}$-gap, which measures the separation between optimal and suboptimal actions at each state.
@@
for a fixed confidence level $\delta\in(0,1)$, typically by combining a union bound with martingale concentration inequalities. While such high-probability guarantees are useful, they only describe the regret distribution at a single point and require the user to choose $\delta$ in advance. By contrast, in many safety-critical applications, such as healthcare decision making, robotic control in human environments, and financial risk management Inamdar et al. (2024), one would like a more detailed characterization of the entire tail distribution of regret, that is, explicit bounds on $\mathbb{P}(R_{K}\geq x)$ that hold for all thresholds $x\geq\mathbb{E}[R_{K}]$.
In this paper, we take a step towards closing this gap by studying the concentration of regret for a UCBVI-type algorithm in finite-horizon MDPs. We show that the tail of the cumulative regret, $\mathbb{P}(R_{K}\geq x)$, admits an explicit, non-asymptotic sub-Gaussian or sub-Weibull bound, depending on the value of $x$ and the instance-dependent $Q^{\star}$-gap parameter.
@@
  • For the $K$-dependent scheme, we show that the regret exhibits a sub-Gaussian tail on an intermediate regime of $x$ and a sub-Weibull tail beyond $x>\tilde{\mathcal{O}}\!\big(H^{5/2}K^{(1+\alpha)/2}\big)$, where $\alpha\in[0,1]$ is a tuning parameter.
  • For the $K$-independent scheme, we obtain a similar two-regime behavior, where the transition to sub-Weibull happens at $\tilde{\mathcal{O}}\!\big(H^{(5-\alpha)/(2-\alpha)}K^{1/(2-\alpha)}\big)$.
  • We further show that the expected regret is $\mathbb{E}[R_{K}]\;\leq\;\tilde{\mathcal{O}}\!\Big(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*}\Big)$, where $\mathrm{gap}^{*}>0$ is the global $Q^{\star}$-gap. The tuning parameter $\alpha$ balances the expected regret and the range over which the regret exhibits a sub-Gaussian tail.
@@ 2 Related Work
The modern theory of online RL in finite MDPs starts from worst-case regret guarantees under the optimism in the face of uncertainty principle. In the average-reward setting, the UCRL2 algorithm of Jaksch et al. (2010) achieves near-minimax regret of order $\widetilde{\mathcal{O}}\!\big(DS\sqrt{AK}\big)$, where $D$ is the diameter of the MDP, and lower bound $\widetilde{\mathcal{O}}\!\big(\sqrt{DSAK}\big)$ shows that the dependence on $(D,S,A,K)$ is almost optimal. Subsequent refinements sharpened the constants and logarithmic factors, for instance by tightening the confidence sets around the transition kernel via empirical Bernstein-type bonuses (Bourel et al., 2020; Fruit et al., 2020).
In the episodic finite-horizon setting considered in this paper, the optimal worst-case scaling was established by the UCBVI algorithm of Azar et al. (2017), which attains regret $\widetilde{\mathcal{O}}\!\big(H\sqrt{SAK}+H\sqrt{KH}\big)$ and derive information-theoretic lower bounds of $\widetilde{\mathcal{O}}\!\big(H\sqrt{SAK}\big)$. A number of works have since refined or extended this result: Efroni et al. (2019) proposed the EULER algorithm, which yields worst-case bounds with improved dependence on the horizon; see also the recent minimax analyses and extensions to factored MDPs, e.g., Tian et al. (2020). These results, however, are primarily worst-case and focus on bounding either the expectation or a single high-probability quantile of the regret, without characterizing the entire tail behavior.
@@
Motivated by the classical gap-dependent theory for stochastic bandits (see, e.g., the survey of Bubeck and Cesa-Bianchi (2012)), there has been a surge of interest in instance-dependent regret bounds for tabular MDPs. Simchowitz and Jamieson (2019) derived the first non-asymptotic gap-dependent regret bounds for model-based tabular RL, showing that the regret can scale polylogarithmically in $K$ (up to polynomial factors in $(S,A,H)$ and $1/\mathrm{gap}^{*}$). Zanette and Brunskill (2019) introduces the environmental norm, a variance-like quantity that controls the difficulty of exploration, and designed an optimistic algorithm whose regret depends on this norm rather than only on worst-case parameters. More recently, Dann et al. (2021) developed improved instance-dependent bounds that replace the global value-function gaps with more local notions tailored to the subset of states visited by optimal policies, together with new information-theoretic lower bounds showing the necessity of these refined gap notions. On the model-free side, Yang et al. (2021) showed that optimistic $Q$-learning enjoys logarithmic regret in episodic tabular MDPs under a global suboptimality-gap on $Q^{\star}$, thus establishing that simple temporal-difference methods can match the best-possible gap-dependent rates of model-based algorithms. Xu et al. (2021) proposed an adaptive multi-step bootstrapping scheme that yields tighter instance-dependent bounds.
Under the linear MDP assumption with function approximation, He et al. (2021) establish instance-dependent logarithmic regret for LSVI-UCB and UCRL-VTR in linear and linear-mixture MDPs. Velegkas et al. (2022) further refine this picture by designing algorithms that incur only a bounded number of policy switches. Most recently, Zhao et al. (2025) extend these guarantees to adversarial environments.
@@
A recent line of work, starting from Fan and Glynn (2024), shows that bandit algorithms that are information-theoretically optimized for Lai–Robbins optimal expected regret can exhibit extremely heavy-tailed regret distributions: the regret tail behaves like a truncated Cauchy, and higher moments grow polynomially in the horizon. Building on this observation, Simchi-Levi et al. (2022, 2023) design policies that explicitly trade off optimality and tail risk, achieving worst-case optimal (or near-optimal) expected regret while simultaneously enforcing exponentially light tails for large-regret events and characterizing sharp Pareto frontiers between expected regret and tail probabilities.
In contrast, RL work that explicitly targets tail or risk-sensitive criteria usually changes the performance objective rather than analyzing the tail of classical regret. For example, there is a growing literature on risk-sensitive or distributional RL that optimizes mean-variance, CVaR, or more general optimized certainty equivalent (OCE) criteria in MDPs; see, e.g., Tamar et al. (2015); Chow et al. (2015); Xu et al. (2023) and references therein. These works derive sample-complexity or regret guarantees for the risk-sensitive value of a policy, not for the tail of the standard regret with respect to the risk-neutral optimal policy.
@@ 3 Problem Formulation
$\mathrm{gap}_{h}(s,a)\;$ $\coloneqq\;\Big(Q_{h}^{*}(s,\pi_{h}^{*}(s))-Q_{h}^{*}(s,a)\Big),\qquad\forall(s,h)\in\mathcal{S}\times[H],$
$\mathrm{gap}^{*}\;$ $\coloneqq\;\min_{\begin{subarray}{c}(s,h)\in\mathcal{S}\times[H]\\ a\notin\mathcal{A}_{h}^{*}(s)\end{subarray}}\mathrm{gap}_{h}(s,a),$ (1)
where $\bm{\pi}^{*}\in\mathbf{\Pi}^{*}$ is any optimal Markov policy.
@@ 4 Main Results
In this section, we present our main results. We study the tail behavior of the regret for a UCBVI-type algorithm in the model-based setting. We analyze a UCBVI-type algorithm that uses the following bonus function:
@@
Here $k\in\{0,\dots,K-1\}$ denotes the episode index, $n\geq 1$ denotes the number of visits to the corresponding state–action pair, and $\alpha\in[0,1]$ and $\mu>0$ are tunable parameters. KD refers to the bonus used in the $K$-dependent algorithm, and KI refers to the bonus used in the $K$-independent algorithm.
The term $K^{\alpha}$ in the definition of $b_{h}^{k}(\cdot)$ plays the role of a global exploration budget that directly controls the aggressiveness of the bonuses. In the original minimax UCBVI analysis of Azar et al. (2017), as well as in the subsequent instance-dependent analysis of Simchowitz and Jamieson (2019), the exploration terms scale only logarithmically with the time horizon, roughly as $\sqrt{{\log(KSAH)}/{n}}$, which is tailored to control of the expected regret. In contrast, our bonuses scale as $\sqrt{(S\ln 2+\mu K^{\alpha})/n}$ (or $\sqrt{(S\ln 2+\mu(k+1)^{\alpha})/n}$ in the $K$-independent case), so that the uncertainty radius around the empirical model grows polynomially in $K$. This more conservative choice incentivizes sustained exploration even in later episodes: when the algorithm has visited a state–action pair only moderately many times, the bonus remains substantial and pushes the optimistic value estimates upward. From a tail perspective, this is precisely what allows us to derive nontrivial bounds on $\mathbb{P}(R_{K}\geq x)$ for large thresholds $x$, at the price of a controlled polynomial dependence on $K$ inside the bonus which will affect the expected regret.
Algorithm 1 Optimistic Value Iteration
1:Input: $H,\mathcal{S},\mathcal{A}$, rewards $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$, initial state $s_{0}$, bonus schedules $\{b^{k}_{h}(\cdot)\}$.
2:Init: For all $(s,a,h)$, $n_{h}^{0}(s,a)=0$ and $\hat{P}_{h}^{0}(\cdot\mid s,a)=\tfrac{1}{S}$, and $\hat{P}_{H-1}^{k}\equiv 0$ and all $k$.
@@
$a_{h}^{k}=\pi_{h}^{k}(s_{h}^{k})\quad\text{and}\quad s_{h+1}^{k}\sim P_{h}^{*}(\cdot\mid s_{h}^{k},a_{h}^{k}),\qquad\forall h\in[H-1],\qquad a_{H-1}^{k}=\pi_{H-1}^{k}(s_{H-1}^{k}).$
6:  Update: For all $(s,a,h)$:
@@
Fix $\alpha\in[0,1]$ and $\mu>0$. For all $\gamma\in[0,1]$, the tail of the cumulative regret $R_{K}$ of Algorithm 1 satisfies
$\mathbb{P}(R_{K}\geq x)\ \leq\ \exp\!\left(-\frac{\big(x-H\lceil K^{\gamma}\rceil-\,m_{K}^{\mathrm{mb}}\big)_{+}^{\,2}}{50H^{5}K}\right)\;+\;\delta_{K}^{\mathrm{mb}}(\gamma),\qquad\forall x\geq H\lceil K^{\gamma}\rceil,$
@@
$m_{K}^{\mathrm{mb}}\;\coloneqq\frac{32H^{6}SA(0.5S\ln 2+\mu K^{\alpha})}{\mathrm{gap}^{*}},$ (2)
@@
$\delta_{K}^{\mathrm{mb}}(\gamma)=SAHK\begin{cases}\displaystyle\exp\left(-2\mu K^{\alpha}\right),&\text{{KD}},\\[8.61108pt] \displaystyle\exp\left(-2\mu K^{\gamma\alpha}\right),&\text{{KI}}.\end{cases}$ (3)
Theorem 1 shows that, under a global $Q^{\star}$-gap, as defined in (1), the tail of the cumulative regret $R_{K}$ admits a decomposition into a sub-Gaussian component and a residual term $\delta_{K}^{\mathrm{mb}}(\gamma)$. The quantity $m_{K}^{\mathrm{mb}}$ is the baseline instance-dependent level of regret around which the distribution concentrates: it aggregates the bonuses incurred before all suboptimal actions have been sufficiently explored and eliminated, and it is of order
$m_{K}^{\mathrm{mb}}\;\approx\;\mathcal{O}\!\big(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*}\big)$
up to logarithmic factors. For thresholds $x$ that are moderately larger than $m_{K}^{\mathrm{mb}}$, the bound $\mathbb{P}(R_{K}\geq x)$ is of sub-Gaussian form with variance proxy $25H^{5}K$, reflecting the Martingale structure of the regret fluctuations after the main exploration phase.
Proposition 1.
@@
$\mathbb{E}[R_{K}]\ \leq\ H\lceil K^{\gamma}\rceil+m_{K}^{\mathrm{mb}}+6KH^{3}\delta_{K}^{\mathrm{mb}}(\gamma),$
where $m_{K}^{\mathrm{mb}}$ is given in (2) and $\delta_{K}^{\mathrm{mb}}(\gamma)$ in (3).
Proposition 1 obtains an instance-dependent upper bound on the expected regret. The first two terms, $m_{K}^{\mathrm{mb}}+H\lceil K^{\gamma}\rceil$, capture the contribution of the baseline instance-dependent level of regret and the initial “burn-in” phase. The final term, proportional to $6KH^{3}\delta_{K}^{\mathrm{mb}}(\gamma)$, accounts for the rare events in which the empirical model falls outside the “good” set at some episodes. In such events the estimated transition kernel $\hat{\mathbb{P}}^{k}$ departs substantially from the true kernel $\mathbb{P}^{*}$ for some $k\in[K]$. Consequently, we cannot ensure that $V^{k}_{0}(s_{0})\geq V^{*}_{0}(s_{0})$ uniformly over $k$, i.e., the optimism property may fail. Thanks to the exponential decay of $\delta_{K}^{\mathrm{mb}}(\gamma)$ in $K^{\alpha}$ (or $K^{\gamma\alpha}$), this residual contribution is negligible: for any fixed $\alpha>0$ and suitable choice of the constant $\mu>0$, the product $K\delta_{K}^{\mathrm{mb}}(\gamma)$ vanishes.
The tuning parameter $\alpha\in[0,1]$ controls the tradeoff between the expected regret and the decay rate of the residual term $\delta_{K}^{\mathrm{mb}}(\gamma)$. Larger values of $\alpha$ make the bonuses more aggressive, which increases the upper bound on the $\mathbb{E}[R_{K}]$ but also causes $\delta_{K}^{\mathrm{mb}}(\gamma)$ to decay faster with $K$ (either as $\exp(-2\mu K^{\alpha})$ or as $\exp(-2\mu K^{\gamma\alpha})$ depending on the bonus schedule). In this sense, $\alpha$ interpolates between prioritizing a smaller expected regret ($\alpha$ small) and enforcing lighter tails for very large deviations ($\alpha$ closer to one). By contrast, the parameter $\gamma\in[0,1]$ is not a tuning knob of the algorithm: it appears only in the analysis and controls the length of the initial burn-in phase $K^{\gamma}$ during which we make no attempt to exploit tail concentration. In Corollary 1, we choose $\gamma$ as a function of $x$ to simplify the resulting tail bound into a clean two-regime description.
@@
$\mathbb{P}(R_{K}\geq x)\leq\begin{cases}\exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{2}}{H^{5}K}\right)\right)&\text{if }\mathcal{O}(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*})\leq x\leq\tilde{\mathcal{O}}(H^{5/2}K^{(1+\alpha)/2})\\ \exp\left(-\tilde{\mathcal{O}}\left(K^{\alpha}\right)\right)&\text{if }x\geq\tilde{\mathcal{O}}(H^{5/2}K^{(1+\alpha)/2})\end{cases}$
@@
$\mathbb{P}(R_{K}\geq x)\leq\begin{cases}\exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{2}}{H^{5}K}\right)\right)&\text{if }\mathcal{O}(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*})\leq x\leq\tilde{\mathcal{O}}(H^{(5-\alpha)/(2-\alpha)}K^{1/(2-\alpha)})\\ \exp\left(-\tilde{\mathcal{O}}\left(\frac{x^{\alpha}}{H^{\alpha}}\right)\right)&\text{if }x\geq\tilde{\mathcal{O}}(H^{(5-\alpha)/(2-\alpha)}K^{1/(2-\alpha)})\end{cases}$
Furthermore, $\mathbb{E}[R_{K}]\leq\mathcal{O}(H^{6}SA(S+K^{\alpha})/\mathrm{gap}^{*})$.